EDF CAC40 World StdDev(%) R Avg return (%)
EDF 1.00 0.88 0.62 17 ?
CAC40 1.00 0.77 13 15
World 1.00 11 12

The above table provides the correlations among the French electric company Electricite de France (EDF), the French stock market index known as the CAC40, and the world market index, together with the standard deviations (SD) of returns and the expected returns R. The risk-free rate is 4.5 percent.

  1. Compute the domestic country beta of EDF as well as its world beta. What do these betas measure?
  2. Suppose the CAC40 stock market is segmented from the rest of the world. Using the CAPM paradigm, estimate the equity cost of capital of EDF.
  3. If EDF decides to issue a Yankee stock offering by cross-listing new shares directly to U.S. public investors on the NYSE, its shares will be tradable internationally. Again, using the CAPM paradigm, estimate EDF’s equity cost of capital under these circumstances.
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

a) Domestic beta=( 17*13*0.88)/13^2= 1.15 World beta= (17*11*0.62)/11^2= 0.958 If the market moves by 1%, the EDF stock retur